OpenStudy (anonymous):

how to solve 2sin(0-47)=1 and find the values?

OpenStudy (zzr0ck3r):

0-47?

OpenStudy (zzr0ck3r):

\(2\sin(\theta-47)=1\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (usman1995):

u kow formula of sine(a-b)? it's sin a x cos b - cos a x sin b,, put vlues in this formula and u will get the answer

OpenStudy (anonymous):

\(\large \rm 2sin(\theta -47)=1 \) \(\large \rm sin(\theta -47)=\frac{1}{2} \) \(\large \rm sin(\theta -47)=\sin\theta \cos47-\cos\theta \sin47=\frac{1}{2} \)